OpenStudy (anonymous):

Solve this equation on the interval 0 θ < 2π. If there is only one solution, write it in the first box and write NONE in the second box. If there is no solution, write NONE in both boxes. Round your answer(s) to two decimal places using radians. -2 cos(θ) + 1 = 2 θ = (smaller value of θ) θ = (larger value of θ)

OpenStudy (anonymous):

the interval is 0( which is less than or equal to) theta ( which is less than) 2pi

OpenStudy (anonymous):

-2cos(x) = 1 cos(x) = -1/2 the related angle is cos inverse of 1/2 , which is 60 degrees

OpenStudy (anonymous):

cos is negative in the 2 and 3 quadrants

OpenStudy (anonymous):

x= 180 -60 and x= 180 + 60 x= 120 , 240 degrees

OpenStudy (anonymous):

\[\frac{2\pi}{3} , \frac{4\pi}{3} \] in radians

OpenStudy (anonymous):

thanks. but the answer seems to be wrong? not sure why

OpenStudy (anonymous):

I am fairly sure what I did was correct

OpenStudy (anonymous):

could it be because the question says round the answer to two decimal points in radians?

OpenStudy (anonymous):

yes lol :|

OpenStudy (anonymous):

obviously if you entered the exact values then they wouldnt like it

OpenStudy (anonymous):

I assumed you would know how to convert the exact values into radians, lol, just calculator work

OpenStudy (anonymous):

yes, however i dont have a calc lol! i think the first one is 2(3.14)/3 ?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=cos%28x%29+%3D-1%2F2+%2C+0%3Cx%3C360 theres a calculator even solves eqns and integrals and everything

OpenStudy (anonymous):

I use it to check my answers

OpenStudy (anonymous):

ok thanks. but to be 100% sure(kind of forgot) radians is using pi as 3.14?

OpenStudy (anonymous):

yes approx that ,

OpenStudy (anonymous):

thanks mate. got it right!

OpenStudy (anonymous):

would you also be able to show me how to do the question just above this question?